Multi-Variable Calculus Assignment

due Tuesday, February 13, 2001

Exercises

1. Find a pair of linear equations that characterize the line given parametrically by

t ⟼ (3, −2, 1) + t[−2, 1, 2] .

2. Let the location of a moving point at time t be given by

(3 cos(t), 3 sin(t), 4t) .

Determine:

(a) The velocity at time t.

(b) The speed at time t.

(c) The acceleration at time t.

(d) The distance traveled by the point between time 1 and time 3.

(e) A pair of equations in x, y, and z that characterize the path traveled.

[Processed from GELLMU to L^{A}TEX on Thu 8 Feb 2001 at 22:14:28 EST]